OpenStudy (richgall____):

I will give you a medal just please help me.. A student writes an incorrect step while checking if the sum of the measures of the two remote interior angles of triangle ABC below is equal to the measure of the exterior angle. Step 1: m∠ m + m∠ n + m∠ o = 180 degrees (sum of angles of a triangle) Step 2: m∠ p – m∠ o = 90 degrees (alternate interior angles) Step 3: Therefore, m∠ m + m∠ n + m∠ o = m∠ o + m∠ p Step 4: So, m∠ m + m∠ n = m∠ pIn which step did the student first make a mistake and how can it be corrected?

OpenStudy (richgall____):

Here Are The Answer Choices Step 1; it should be m∠m + m∠ n + m∠ o = 90 degrees (corresponding angles) Step 1; it should be m∠m + m∠ n + m∠ o = 90 degrees (adjacent angles) Step 2; it should be m∠o + m∠ p = 180 degrees (alternate exterior angles) Step 2; it should be m∠o + m∠ p = 180 degrees (supplementary angles)

OpenStudy (richgall____):

Anyone?

OpenStudy (richgall____):

@Vivicat

OpenStudy (anonymous):

Step 2; it should be m∠o + m∠ p = 180 degrees (alternate exterior angles) << I think this is it

OpenStudy (anonymous):

what do yo think?

OpenStudy (richgall____):

I think it is too